OpenStudy (anonymous):

Rationalize the denominator 3 divided by sqrt2?

OpenStudy (anonymous):

is it 3sqrt2/2

OpenStudy (anonymous):

gotta multiply by that "clever form of one"\[\frac{3}{\sqrt{2}}\cdot 1=\frac{3}{\sqrt{2}}\cdot \frac{\sqrt{2}}{\sqrt{2}}=\frac{3\sqrt{2}}{2}\]

OpenStudy (anonymous):

Thanx